OpenStudy (anonymous):

A marketing group noticed a trend in running shoes sales after they introduced a new product into the market. # of Days # of Pairs of Shoes sold (thousands) 7 0.3 14 0.7 21 1.9 28 3.2 35 5.0 Which of the following is the Exponential Regression equation that could be used to determine the number of pairs of shoes sold given the number of days after introduction?

OpenStudy (anonymous):

y = 0.17(1.107)^x y = 0.3(0.6)^x y = 1.7(0.3)^x y = 0.3(6)^x

OpenStudy (anonymous):

@mertsj can u help me